OpenStudy (anonymous):

12. Suppose you did a survey of male shoes sizes and the mean of that survey was size 10.5 with a standard deviation of 1.5. What is the z-score of a size of 11? A. 0.75 B. 0.50 C. 0.67 D. 0.33

OpenStudy (anonymous):

??

OpenStudy (campbell_st):

well its simply \[z = \frac{x - \mu}{\sigma}\] so in your question \[z = \frac{11 - 10.5}{1.5}\] just calculate it out

OpenStudy (campbell_st):

the z score formula is \[z = \frac{score - mean}{standard..deviation}\] hope this helps

OpenStudy (anonymous):

I got 4 , for my answer

OpenStudy (campbell_st):

try it this way \[z = \frac{(11 - 10.5)}{1.5}\] as the answer isn't 4